OpenStudy (anonymous):

mark mcguire hits a baseball so that it leaves that bat at a speed of 30 m/s and at an angle of 36.9 degree above the horizontal. ignore air resistance.(a).what two times is the baseball at a height of 10 cm above the point at which it left the bat?(b).calculate the horizontal and vertical components of baseballs velocity at each of the two times calculated in part (a) (c).what is the magnitude and direction of the baseballs velocity when it returns to the level at which it left the bat?

OpenStudy (anonymous):

so you are solving for deltay so use the equation \[\Delta y=(v _{0}\sin \theta _{0} )t-\frac{ 1 }{ 2 } g(t)^2 \]

OpenStudy (anonymous):

you have the angle, the initial speed and Delta y because you are looking for 10 g is gravity acceleration

OpenStudy (anonymous):

so \[10=(30*\sin36.9)t-\frac{ 1 }{ 2 }*9.81*t^2\]

OpenStudy (anonymous):

solve for t

OpenStudy (anonymous):

thanks